in a club, the ratio of boys to girls is 7:8. the ratio of one less than the number of girls to the total number of members is 1:2. how many girls are in the club?
Answer by mananth(16949)

boys/girls = let girls be x
boys be y
y/x= 7/8
y= 7x/8 ................1
..
x-1 / x+y = 1/2
2(x-1)=x+y
2x-2=x+y
2x-x-y=2
x-y =2...................2
plug value of y in (2)
x-7x/8 =2
8x-7x / 8 = 2
x/8 =2
x=16 girls
...
m.ananth@hotmail.ca